Determine if the following are in proportion. 
15, 45, 40, 120 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

15, 45, 40, 120
15 / 45 = 1 / 3
40 / 120 = 1 / 3
Hence, 15: 45 = 40:120
∴These are in a proportion

Thanks